Amy is making a pitcher of juice for her party. She has two different recipes to choose from. Welch's Grape Juice = 3 cups grape concentrate, 5 cups cold water. The other is Brand X Grape Juice = 2 cups grape concentrate, 3 cups cold water.
What fraction of each recipe is concentrate? 3/8, & 2/5
What percent of each recipe is concentrate? 3/8=37.5=38 percent, & 2/5=40 percent
Which recipe should Amy choose if she wnts her juice to have the most grape flavor? the Brand X Grape Juice
Is this right?

Answer by Theo(13342)

welch = 3 parts grape and 5 parts water
other = 2 parts grape and 3 parts water
welch grape fraction is 3/8
other grape fraction is 2/5

welch grape percent is 3/8 = .375 * 100% = 37l5%
other grape percent is 2/5 = .4 = 40%

for more grape flavor, she should use other.

you are correct.